# On the Power Saving in High-Speed Ethernet-based Networks for Supercomputers and Data Centers

### v2

@@ 2.5 Proposals to complement EEE
The main basis behind power management in interconnection networks is that link power consumption does not depend on usage. However, link usage is not constant because transmissions only occur when necessary. Moreover, the usage of any particular link is determined by several factors. For example, the traffic pattern, the topology, and the routing algorithm can have different effects on the usage of every link during the execution of an application. Deciding on when to switch to and from a power-saving state is important because the more power savings, the more overhead the link suffers.
@@
One option to reduce transitions is frame buffering: accumulating packets on buffers and only transitioning to Wake on certain conditions [christensen_ieee_2010, saravanan_powerperformance_2013]. Once a packet is on an output port buffer or is granted crossbar traversal on switches, an occupation threshold and a time limit are set. The buffer occupation is set to exploit packet coalescing and thus, avoid transitioning to Wake once per packet. However, a timeout is needed to ensure packets are eventually sent regardless of the buffer occupation. Obviously, this can cause additional latency for every hop. In the worst-case scenario, every communication is composed of small and spaced out packets that do not coincide in time or share network paths. This means that packets will be stalled at every port and have their latency increased both by the buffering timeout and the transition time to Wake. Of course, this is not viable for latency-sensitive applications. As we cannot assume packet coalescing to mitigate the overhead of port transitions, the only option left is to keep links in the Wake state for longer. If the port is left in Wake state for a moderate period after transmission, there is no overhead to a packet that must be transmitted within that time. Power Down Threshold (PDT) [saravanan_powerperformance_2013] was proposed so that every port stays operational for an additional time after transmission, at the cost of potential power saving.
A simple implementation of PDT can be done by using a timer on each port. The port switches to power-saving mode when the timer expires without transmitting a packet. If any packet arrives at the port, either from the link or via arbitration or fragmentation, the timer is canceled. It will be set again on the last port that has transmitted through the link, once it has no packets to send.
Figure 4 shows link behavior when using PDT in terms of power consumption over time. The blue arrows represent the duration of the timer used in PDT, and the colors of the timer icon represent changes in the timer used in PDT. The green-colored timer means that it is set, while the red indicates that the timer is canceled, and the salmon is for the timer expiration. We can see that during the Wake state, packets present at the port can be sent normally. After sending a packet, the port waits for a certain period (${t}_{PDT}$) to transition to the Sleep state. If a packet arrives within ${t}_{PDT}$ after sending a packet, the timer is canceled and set again when the packet is transmitted.
@@
This way, ports that may frequently transmit will be in the Wake state most of the time, but packets will not suffer from latency because ports are ready to send them. Meanwhile, for inactivity periods that are on the order of seconds, the overhead on latency for state transitioning is irrelevant. A ${t}_{PDT}$ value of microseconds is also insignificant compared to the inactivity period in terms of power saving. The issue, as we have mentioned, is the intermittent nature of network activity during communication. When the ${t}_{PDT}$ is on the same order of magnitude as the inactivity periods, ports may often receive packets after powering down, or even while transitioning.
@@
The main inconvenience of the PDT technique is the criteria for choosing a value and its generalization to all network ports. The former depends on the application and would require prior execution for inactivity profiling. This would be useful if the inactivity periods of the ports could be estimated beforehand. However, they depend on multiple factors, such as task mapping and the number of processors on the nodes, which affect the traffic pattern. In a production environment, the values would have to be set depending on the application. The latter issue of PDT is that inactivity periods differ between all ports. This can be caused by network topology or task scheduling. The topology determines the average and maximum hop count for packets between a source and destination, both by the layout of network elements and the routing algorithm. In fact, the higher the hop count, the larger the overhead due to port transitions.
@@
Due to these limitations, ports should establish their own ${t}_{PDT}$ value, based on previous network activity. To solve this issue, Saravanan proposed PerfBound [perfbound]. The basis of PerfBound is forming a histogram on each port with the inactivity periods so that port can calculate its particular ${t}_{PDT}$ at any given moment in a way that degradation is bound. PerfBound also requires every port to store (or calculate) the distance in hops of any packet, given a destination. It uses one counter per possible distance value to calculate the average distance of previous packets. In this way, we can quantify the possible degradation of a packet due to the powering up of multiple ports along its route.
@@
To calculate the next ${t}_{PDT}$ value, the algorithm first determines the number of packets that can be delayed due to power-up transitions within a given degradation limit. For example, if we want to maintain a performance degradation restriction of up to 1 % in a period of duration $X$, we can delay up to $N=0.01*X/{t}_{w}$ packets, where ${t}_{w}$ is the transition time to Wake. We use the average distance to destinations described earlier as a corrective factor $l$ for each port. By doing so, we are less restrictive on ports with higher mean distance. In a generalized manner, $l=(bound\%)\times\sum_{i=1}^{n}\frac{p_{i}}{h_{i}}$, where $n$ is the number of different hop counts, $p_{i}$ is the proportion of packets that travel $h_{i}$ hops and $h_{i}$ is the number of hops for the $i$-th group of packets.
@@
For example, if 60 % of packets sent by a port are 6 hops away from the destination, and the distance for the remaining 40 % is 6 hops, we would calculate that factor for a 1 % degradation threshold as in Eq. 1:
$l=0.6\times\frac{0.01}{4}+0.4\times\frac{0.01}{6}=0.01\times\left(\frac{0.6}{4}+\frac{0.4}{6}\right)\approx 0.0022$ (1)
@@
To apply this factor, the new number of packets will be $N=l*X/{t}_{w}$. Once we have obtained $N$, we search on the histogram, from highest to lowest bin, and accumulate the count of each bin. We are looking for the leftmost bin that accumulates at most $N$. We will set the ${t}_{PDT}$ to the mean of that bin.
@@ 2.6 BXIv3 network technology
BXIv3 [bxi3eviden] is the network technology being developed by Eviden for HPC systems and data centers. The main objective for Eviden is to develop an interconnect that can power post-exascale systems. The existing versions of BXI, namely V1.2 [bxihoti] and V2 [bxi2atos], are already deployed in several TOP500 supercomputers, including CEA-HF and Tera-1000-2 [top500]. Choosing Ethernet as a basis means that there is a grounded work on standards that can speed up the development of the technology. Some examples of this set of standards are the 802.1Qbb or priority flow control (PFC) [cclsrdmad] for lossless per-queue link-level flow control, or the possibility to use variable-sized packets and jumbo frames.
BXIv3 network interfaces (NICs) support a link bandwidth of $400$ Gbps (expected to be $800$ Gbps in the near future), with a target deployment scale of up to 100,000 nodes, following the modular supercomputing architecture (MSA) specification [Suarez19MSA]. BXIv3 supports different network topologies, such as Fat-trees and Megafly/Dragonfly+ [flajslik_megafly_2018, dfly+], which permit the desired levels of interconnection scalability. BXIv3 switches assume a Combined Input-Output Queued (CIOQ) switch architecture, similar to other Ethernet-based options. Regarding power consumption, we are not aware of any proposal tailored to BXIv3222Further details on the BXIv3 architecture can be obtained from one of the papers describing the BXIv3 architecture in the RED-SEA project [Gomez24redsea].. We believe that, given the high consideration for power consumption, it will be a cornerstone for future interconnection technologies.
@@ 3 Power consumption model
This section describes the interconnection network power model tailored to the BXIv3 architecture we implemented in our network simulator. This model is based on the Energy Efficient Ethernet (EEE) standard (see Section 2.4 and the extensions proposed to expand its functionality (see Section 2.5), such as LPI, PDT, and Perfbound. We have also designed some improvements to the Perfbound strategy, so that EEE can be better applied to Ethernet-based interconnection networks when used to run HPC applications.
The EEE standard specifies the PHY layer behavior, but the transition between link states is left to the vendor. For BXIv3-based interconnection networks, we assume a combined input-output queued (CIOQ) switch architecture. Specifically, BXIv3-based switches define buffers at input and output ports, so it is possible to grant switch crossing requests from input buffers to non-operative ports, i.e, to output ports in the Sleep state. This way, packets will be stored on their respective output queues when ports power up, i.e., when returning to the Wake state.
@@ 3.1 Link power-levels and transition logic
The link power levels are as described in the EEE standard, with three states: Wake, Fast Wake, and Deep Sleep, and as we showed in Table 3. As the EEE only specifies the components to shut down and the corresponding timings to transit between them, it is also up to the vendors and network architects to decide when and how to switch states.
@@ 3.2 PerfBound implementation details
For example, the histogram formed of inactivity period lengths is the key piece of the mechanism. However, even if clearing the histogram (or refreshing, as the authors mention) is pondered, its consequences are only theoretically mentioned. There may be several possibilities for managing the histogram to constrain its storage size and maintain a bias towards recent values.
Another important aspect of the algorithm is using average hop distances from switches and nodes to destination nodes. The experiments exploring the original work set a constant hop count for the network. However, this is non-realistic and may be affected by the traffic patterns of different applications. Devices should have their own record of hop distances, based on the destinations that can be reached according to their routing algorithm. Distance calculations depend on the topology and routing employed, and should be straightforward on regular topologies. On irregular topologies, however, hop distances should be calculated beforehand. Employing routing information may also decrease the effectivity of the algorithm if the routing algorithm is adaptive and hop counts change depending on the network status.
@@
The recurrent calculation of a PDT timer expiration value for every port is also concerning. More so for switches, which would need to do so every time all output queues for any port are emptied. We believe that this processing should be carried out on a regular basis, but only if there was network activity on a port since the last calculation.
@@ 3.3 Additional requirements
All network equipment already has the circuitry and mechanisms to implement the different power-saving states defined in the EEE standard. However, vendors must define the logic to switch between port power states.
Using PDT requires, per port, one timer and a separate register to store the configured ${t}_{PDT}$. The time resolution and maximum possible value are determining factors. For instance, a 32-bit unsigned register with up to nanosecond precision can encode ${2}^{32}*\frac{1s}{10^{9}ns}\approx 4.295$ seconds. Table 4 shows the maximum time that could be encoded in a register with several register widths and various time resolutions.
@@
Width
[…]
32 bits $4.3\,\mathrm{s}$ $43\,\mathrm{s}$ $429\,\mathrm{s}$ $4\times 10^{3}\,\mathrm{s}$
48 bits $2.81\times 10^{5}\,\mathrm{s}$ $3\times 10^{6}\,\mathrm{s}$ $28\times 10^{6}\,\mathrm{s}$ $2.81\times 10^{8}\,\mathrm{s}$
64 bits $1.8\times 10^{10}\,\mathrm{s}$ $1.84\times 10^{11}\,\mathrm{s}$ $2\times 10^{12}\,\mathrm{s}$ $1.8\times 10^{13}\,\mathrm{s}$
128 bits $3.4\times 10^{29}\,\mathrm{s}$ $3\times 10^{30}\,\mathrm{s}$ $3.4\times 10^{31}\,\mathrm{s}$ $3.4\times 10^{32}\,\mathrm{s}$
@@
For PerfBound, each port requires storage to form the histogram. The most naive implementation would be storing each inactivity duration in a sorted list to facilitate the calculi of the next ${t}_{PDT}$ value. However, it would be more feasible to set a constant bin width and store bin counts instead, at the cost of precision. An improvement on this would be to use a map-based structure instead. This way, recurrent bins do not increase storage size. As we are storing bin indices instead of raw time differences, the width does not have to coincide with that of the ${t}_{PDT}$.
With finite memory, regular clean-up of the recorded inactive periods should be carried on. If the histogram is cleared every $N$ elements collected, a possible storage solution is a FIFO queue with at most $N$ bin indices. Then, the histogram can be formed by the elements in the queue. Alternatively, only the oldest element can be removed if the histogram is cleared periodically without any other concern, and the size is indefinite. However, steep memory requirement increases are not expected if cleared within a sensible period, like one second, at most. Moreover, the packet rate and size to cause this would be extremely inefficient because a significant part of the traffic would be the overload from packet headers. And even then, those inactive periods would fall under a few bins. Ultimately, the space required to store inactivity periods is determined by bin width (${w}_{bin}$) and the highest possible bin (${max}_{bin}$), with up to $\max_{\text{bin}}/w_{\text{bin}}$ total counts. If the histogram is cleared every ${TTL}_{hist}$ seconds, it cannot have more than $\text{TTL}_{\text{hist}}/w_{\text{bin}}$ different bins. If the decision to clear the histogram is based on the number of values collected, the largest histogram would be one of $N$ bins with count 1.
@@ 3.4 Improvements to Perfbound: PerfBoundCorrect
Even if PerfBound is a grounded basis for predicting a reasonable ${t}_{PDT}$ value, it presents mainly two drawbacks. The first one is that collecting previous values to calculate a new ${t}_{PDT}$ is ultimately a prediction of a concrete value. Prediction is widely used in decision-making on computers, from branching predictors to unattended maintenance at low-impact intervals.
However, missing predictions often have no more performance penalties than the overheads obtained by not using the predictor. But this is not the case for PerfBound, and the performance impact of every missed prediction is quantifiable. We can calculate the ratio between a period of inactivity that caused a port to power down and the last ${t}_{PDT}$ value. That proportion can help quantify the error factor in missed predictions. If we collect several of these values, we can calculate a corrective factor to increase the ${t}_{PDT}$ value according to the missing rate and ratios. We call this technique PerfBoundCorrect.
@@
We need a shift register and a queue containing the last ratios to track the last predictions. Every time a packet arrives at a port, if the PDT is still set (the port is up), we count this as a hit, or a miss otherwise. We can store this history in a shift register of length $n_{R}$. When a port starts its transition to power up, the ratio is calculated and stored in a FIFO queue of length $n_{R}$ at the same time as the shift register is updated. If we remove a miss from the register by shifting, we also have to pop the oldest element from the queue. This way, we keep as many elements on the queue as misses encoded at the shift register. The proportion between the miss count and the shift register length returns the hit rate of the last $n_{R}$ predictions, and the queue of ratios can be used to get a correcting factor. It is important to note that a greater ratio of outliers is expected and reasonable, since they appear when a long inactivity period has finished. In those cases, the overhead of transitions is acceptable because they are the best opportunities for power saving.
@@
To apply the correcting factor to the PerfBound calculation, we have opted for the geometric mean of all ratios times the miss rate for the last $T_{PDT}$ value. The geometric mean was chosen because it is more tolerant of outliers than the arithmetic mean. The corrective factor ($cf$) applied to the $t_{PDT}$ calculation is as follows:
@@
$cf=miss\%\times\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=miss\%\times{\sqrt[n]{r_{1}r_{2}\cdots r_{n}}}$
@@
To prevent too small $cf$ values and result in unreasonably high (i.e., second range or above) $t_{PDT}$ values, we set a maximum limit $max_{t_{PDT}}$ within a sensible order of magnitude. The corrective factor $cf_{b}$ will be $min(cf,max_{t_{PDT}})$. Giving this leeway to the PerfBound algorithm may seem harmful to power saving as links would spend more time operational after their transmissions end. However, it also has the potential to reduce the overhead caused by very reduced ${t}_{PDT}$ values. This is our proposal, and the results are non-trivial, as we will see in Section 4.

### v3

@@ 2.4 Proposals to complement EEE
The main basis behind power management in interconnection networks is that link power consumption does not depend on usage. However, link usage is not constant because transmissions only occur when necessary. Moreover, the usage of any particular link is determined by several factors. For example, the traffic pattern, the topology, and the routing algorithm can have different effects on the usage of every link during the execution of an application. Deciding on when to switch to and from a power-saving state is important because the more power savings, the more overhead packets can suffer.
@@
The overhead for turning a port on is greatly compensated if the following transmission is comparatively higher. However, short and sparse packets increase transmission overhead, similarly. If the port is left in Wake state for a moderate period after transmission, there is no overhead to the next packet transmitted within that time. Power-down Threshold (PDT) [30] was proposed so that every port stays operational for an additional time after transmission, at the cost of potential power saving.
The technique can be implemented with a timer on each port. The port turns down when the timer expires after transmitting a packet. If any packet arrives at the port, either from the link or via arbitration or fragmentation, the timer is canceled. It is set again right after transmission, when there are no more packets to send.
Figure 4 shows link power behavior under PDT. Blue arrows indicate the timer duration, and the timer icon’s color reflects its state changes. Green means that it is set, red is for reset, and salmon is for expired. During the Wake state, packets present at the port can be sent normally. After sending a packet, the port waits for a certain period (${t}_{PDT}$) to transition to the Sleep state. If a packet arrives within ${t}_{PDT}$ after sending a packet, the timer is canceled and set again when the packet is transmitted.
@@
This way, frequently used ports will be in the Wake state most of the time, but packets will not suffer from latency because ports are ready to send them. Meanwhile, for inactivity periods that are on the order of seconds, the overhead on latency for state transitioning is irrelevant. The issue, as we have mentioned, is the intermittent nature of network activity during communication. When the ${t}_{PDT}$ is on the same order of magnitude as the inactivity periods, ports may often receive packets after powering down, or even while transitioning.
@@
The main drawback of PDT is selecting an appropriate timer value and generalizing it for all ports. The former depends on the application and would require prior execution for inactivity profiling. This would be useful if the inactivity periods of the ports could be estimated beforehand. However, they depend on multiple factors, such as task mapping and the number of processors on the nodes, which affect the traffic pattern. In a production environment, the values would have to be fine-tuned for an application. The latter issue of PDT is that inactivity periods differ between all ports. This can be caused by network topology or task scheduling. The topology determines the average and maximum hop count for packets between a source and destination, both by the layout of network elements and the routing algorithm. In fact, the higher the hop count, the larger the overhead due to port transitions.
@@
Due to these limitations, ports should establish their own ${t}_{PDT}$ value, based on previous network activity. To solve this issue, Saravanan proposed PerfBound [31]. The basis of PerfBound is forming a histogram on each port with the inactivity periods so that the port can calculate its particular ${t}_{PDT}$ at any given moment in a way that degradation is bound. PerfBound also requires every port to store (or calculate) the distance in hops of any packet, given a destination. In this way, we can quantify the possible degradation of a packet due to the powering up of multiple ports along its route.
@@
To calculate the next ${t}_{PDT}$ value, the algorithm first determines the number of packets that can be delayed due to power-up transitions within a given degradation limit. For example, if we want to maintain a performance degradation restriction of up to 1 % in a period of duration $X$, we can delay up to $N=0.01*X/{t}_{w}$ packets, where ${t}_{w}$ is the transition time to Wake. We use the average distance to destinations described earlier as a corrective factor $l$ for each port. By doing so, we are less restrictive on ports with a higher mean distance. In a generalized manner, $l=(bound\%)\times\sum_{i=1}^{n}\frac{p_{i}}{h_{i}}$, where $n$ is the number of different hop counts, $p_{i}$ is the proportion of packets that travel $h_{i}$ hops and $h_{i}$ is the number of hops for the $i$-th group of packets.
@@
For example, if 70 % of packets sent by a port are 6 hops away from the destination, and the distance for the remaining 30 % is 6 hops, we would calculate that factor for a 1 % degradation threshold as in Eq. 1:
$l=0.7\times\frac{0.01}{4}+0.3\times\frac{0.01}{6}=0.01\times\left(\frac{0.6}{4}+\frac{0.4}{6}\right)\approx 0.0023$ (1)
@@
To apply this factor, the number of packets we can delay will be $N=l*X/{t}_{w}$. Once we have obtained $N$, we search on the histogram, from the highest to the lowest bin, and accumulate the count of each bin. We are looking for the leftmost bin that accumulates at most $N$. We will set the ${t}_{PDT}$ to the mean of that bin.
@@ 2.5 BXIv3 network technology
BXIv3 [15] is the network technology being developed by Eviden for HPC systems and data centers. The main objective for Eviden is to develop an interconnect that can power post-exascale systems. The existing versions of BXI, namely V1.2 [12] and V2 [6], are already deployed in several TOP500 supercomputers, including CEA-HF and Tera-1000-2 [38]. Choosing Ethernet as a basis means that there is a grounded work on standards that can speed up the development of the technology. Some examples of this set of standards are the 802.1Qbb or priority flow control (PFC) [42] for lossless per-queue link-level flow control, or the possibility to use variable-sized packets and jumbo frames.
BXIv3 network interfaces (NICs) support a link bandwidth of $400$ Gbps (expected to be $800$ Gbps in the near future), with a target deployment scale of up to 100,000 nodes, following the modular supercomputing architecture (MSA) specification [36]. BXIv3 supports different network topologies, such as Fat-trees and Megafly/Dragonfly+ [16, 34], which permit the desired levels of interconnection scalability. BXIv3 switches assume a Combined Input-Output Queued (CIOQ) switch architecture, similar to other Ethernet-based options. Regarding power consumption, we are not aware of any proposal tailored to BXIv3222Further details on the BXIv3 architecture can be obtained from one of the papers describing the BXIv3 architecture in the RED-SEA project [14].. We believe that, given the high consideration for power consumption, it will be a cornerstone for future interconnection technologies.
@@ 3 Power consumption model
This section describes the interconnection network power model tailored to the BXIv3 architecture we implemented in our network simulator. This model is based on the Energy Efficient Ethernet (EEE) standard (see Section 2.3) and the extensions proposed to expand its functionality (see Section 2.4), such as LPI, PDT, and PerfBound. We have also designed some improvements to the PerfBound strategy, so that EEE can be better applied to Ethernet-based interconnection networks when used to run HPC applications.
For BXIv3-based interconnection networks, we assume a combined input-output queued (CIOQ) switch architecture. Specifically, BXIv3-based switches define buffers at input and output ports, so it is possible to grant switch crossing requests from input buffers to non-operative ports, i.e, to output ports in the Sleep state. This way, packets will be stored on their respective output queues when ports power up, i.e., when returning to the Wake state.
@@ 3.1 Link power-levels and transition logic
The link power levels are as described in the EEE standard, with three states: Wake, Fast Wake, and Deep Sleep. As the EEE only specifies the components to shut down and the corresponding timings to transit between them, it is also up to the vendors and network architects to decide when and how to switch states.
@@ 3.2 PerfBound implementation details
For example, the histogram formed of inactivity period lengths is the key piece of the mechanism. However, even if clearing the histogram (or refreshing, as the authors mention) is pondered, its consequences are only theoretically mentioned. There may be several possibilities for managing the histogram to constrain its storage size and maintain a bias towards recent values. Deciding on which information to remove from the histogram, based on time or size factors, can lead to different results, but it is necessary as storage is finite. The maintenance parameters of the histograms are shown in Table 4.
Parameter Description Encodes
$n_{vals}$ # of stored values Counter
${n}_{bins}$ # of populated histogram bins Counter
$r_{bin}$ Bin/element range (time unit) Time
$w_{counter}$ Bin counter width Size (bytes)
$s_{elem}$ Element size Size (bytes)
${TTL}_{hist}$ Histogram maximum time-to-live Time
$h_{maxN}$ Maximum records on histogram Counter
$h_{maxval}$ Maximum value to store in the histogram Time
Table 4: Histogram maintenance parameters
The behavior upon collecting an inactivity time with value $t_{inac}$ (up to $h_{maxval}$) depends on the histogram management strategy. If the criterion to clean the histogram is to clear the whole structure, we only have to find the bin index for $t_{inac}$ and store the bin counter. $t_{inac}$ falls into bin ${inac}_{i}/r_{bin}$, where $r_{bin}$ represents a time unit, and each bin indexes a counter of $w_{counter}$ bytes. A map-like structure is ideal for this implementation, only storing counters for populated bins.
When the timer to wipe the histogram expires after ${TTL}_{hist}$ or after storing $h_{maxN}$ elements, the map is cleared. If we want to update the histogram in an informed manner, for example, by getting rid of the oldest values, we must track individual values and construct the histogram using them. We would implement this, for example, with a circular buffer of $h_{maxN}$ elements. Using that list, we can calculate the bin indices for each element and increment bin counters. The space requirements of these histogram implementations will be further described in section 3.3. In any of those cases, we can also set a ${haxval}_{hist}$ to prevent great outliers and reduce the total number of bins in the histogram.
Another important aspect of the algorithm is using average hop distances from switches and nodes to destination nodes. The experiments exploring the original work set a constant hop count for the network. However, this is non-realistic and may be affected by the traffic patterns of different applications. Devices should have their own record of hop distances, based on the destinations that can be reached according to their routing algorithm. Device-to-destination distance can either be calculated theoretically (offline) using the routing algorithm or dynamically obtained with a traceroute-like all-to-all operation. For the second option, every source node sends this kind of message to each possible destination, increasing the TTL until it reaches the real hop count. Any intermediary device that sets TTL to 0 will then record the highest TTL value that subsequent messages from the same source to the same destination have. The highest TTL value recorded by a device will contain the exact hop count to the destination.
@@
The recurrent calculation of a PDT timer expiration value for every port is also concerning. More so for switches, which would need to do so every time all output queues for any port are emptied. We believe that this processing should be carried out regularly, but only if there was network activity on a port since the last calculation.
@@ 3.3 Additional requirements
PDT requires one timer and one register to store the configured ${t}_{PDT}$ in each port. The time resolution and maximum possible value are determining factors. For instance, a 32-bit unsigned register with up to nanosecond precision can encode ${2}^{32}*\frac{1s}{10^{9}ns}\approx 4.295$ seconds. Table 5 shows the maximum time that could be encoded with several register widths and time resolutions.
@@
Width 16 bits 65.54 µs 655.36 µs 6.554 ms 66 ms
32 bits 4.295 s 43 s 429 s $4\times 10^{3}$ s
48 bits $2.81\times 10^{5}$ s $3\times 10^{6}$ s $28\times 10^{6}$ s $2.81\times 10^{8}$ s
64 bits $1.8\times 10^{10}$ s $1.84\times 10^{11}$ s $2\times 10^{12}$ s $1.8\times 10^{13}$ s
128 bits $3.4\times 10^{29}$ s $3\times 10^{30}$ s $3.4\times 10^{31}$ s $3.4\times 10^{32}$ s
@@
For PerfBound, each port requires some memory to form the histogram. For an estimation of the histogram memory footprint, we must look again at Table 4. Mainly, the maximum histogram size depends on its maintenance strategy and the width for the bin counters ($w_{counter}$). The inactivity periods will be determined by the traffic pattern of the application, which will generate a variable number of histogram bins. If we only store the counters for every histogram bin, the histogram will take up $w_{counter}\times{n}_{bins}$ bytes, with ${n}_{bins}\leq h_{maxN}$. For a histogram of up to $h_{maxN}$ elements stored in a circular buffer, the extra memory requirement is $h_{maxN}\times s_{elem}$. In other words, up to $h_{maxN}*(w_{counter}+s_{elem})$ bytes of memory if ${n}_{bins}=h_{maxN}$. Steep memory requirement increases beyond the range of MiBs are not expected if the histogram is cleared within a sensible period, like one second, and values are below the range of millions, with the original authors pointing out 20k elements as an example.
@@ 3.4 Improvements to PerfBound: PerfBoundCorrect
Even if PerfBound is a grounded basis for predicting a reasonable ${t}_{PDT}$ value, it presents mainly two drawbacks. The first one is that collecting previous values to calculate a new ${t}_{PDT}$ is ultimately a prediction for the next one. Prediction is widely used in decision-making on computers, from branching predictors to unattended maintenance at low-impact intervals. In several of these scenarios, missing predictions often have the same performance overhead as not using the predictor; however, for PerfBound, the performance impact of every missed prediction is quantifiable.
Every time we turn on a link (after $t_{inac}$ seconds of inactivity), it means that our last prediction missed because the last $t_{PDT}$ value is lower than $t_{inac}$. In other words, $t_{PDT}$ should have been ${t_{inac}}\over{t_{PDT}}$ (>1>1) times longer for the prediction to hit. By measuring the frequency and magnitude of misses on recent predictions, we aim to enhance PerfBound with a correction factor that reduces future miss rate and thus, the performance overhead. We call this technique PerfBoundCorrect. The only new parameter for this enhancement is the history length.
As this technique is built upon PerfBound, the same space requirements and histogram management strategies apply. Our proposal employs the output of the PerfBound technique and previous predictions to calculate a new $t_{PDT}$ value. The sequence is: collect inactivity periods, form the history/histogram, calculate the PerfBound prediction, and then calculate the PerfBoundCorrect factor. The enhancements to power efficiency come from applying the corrective factor, as the histogram is managed similarly.
Memory-wise, we have to store the miss history of the previous $n_{R}$ predictions, and the ratios that are above 1.
@@
The miss history can be encoded in a shift register, and a FIFO queue of up to $n_{R}$ floating-point numbers can be used for the ratios. Every time a packet arrives at a port, if the PDT is still set (the port is up), we count this as a hit, or a miss otherwise. When a port starts its transition to power up, the ratio (${t_{inac}}/{t_{PDT}}$) is calculated and stored in a FIFO queue of length $n_{R}$ at the same time as the shift register is updated. If we remove a miss from the register by shifting, we also have to pop the oldest element from the queue. This way, we keep as many elements on the queue as misses encoded at the shift register. We only track missed predictions because we want to exclusively increase $t_{PDT}$ values using those ratios above 1. The proportion between the miss count and the shift register length returns the hit rate of the last $n_{R}$ predictions, and the queue of ratios can be used to calculate a correcting factor for the output of PerfBound.
@@
To apply the correcting factor to the PerfBound calculation, we have opted for the geometric mean of all ratios times the miss rate for the last $T_{PDT}$ value. The geometric mean was chosen because it is more tolerant of outliers than the arithmetic mean. The corrective factor ($cf$) applied to the output of PerfBound is as in Eq. 2:
@@
$cf=miss\%\times\left(\prod_{i=1}^{n}x_{i}\right)^{\frac{1}{n}}=miss\%\times{\sqrt[n]{r_{1}r_{2}\cdots r_{n}}}$ (2)
@@
To prevent too small $cf$ values and result in unreasonably high (i.e., second range or above) $t_{PDT}$ values, we set a maximum limit $h_{maxval}$ within a sensible order of magnitude. The corrective factor $cf_{b}$ will be $min(cf,h_{maxval})$. The impact of great outliers in ratios that come from long inactivity periods is minimized by the miss rate: the lower the miss rate, the less these values affect the correcting factor. The influence of these outliers in the corrective factor diminishes as the network activity increases over time, and we reduce the miss rate. Giving this leeway to the PerfBound algorithm may seem harmful to power saving, as links would spend more time operational after their transmissions end. However, it also has the potential to reduce the overhead caused by very reduced ${t}_{PDT}$ values. This is our proposal, and the results are non-trivial, as we will see in Section 4.